OpenStudy (anonymous):

Given R= xi + yj + zk, and r=l R l, determine:

OpenStudy (anonymous):

(a)\[\nabla r ^{n}\] (b)\[\nabla . (r ^{n} R)\]

OpenStudy (amistre64):

if I could read the notation i might have a clue :/ gradient r^n I got no idea what that even implies

OpenStudy (anonymous):

lol i tried reading up materials from ka stroud, james, etc etc and i couldnt understand the question also.. T_T

OpenStudy (amistre64):

im thinking it might be russian :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=grad+r%5En

OpenStudy (amistre64):

since r = sqrt(x^2+y^2+z^2) hmmm

OpenStudy (amistre64):

\[r^n = (x^2+y^2+z^2)^{n/2}\] but i got no clue what thats spose to be

OpenStudy (amistre64):

unless, we take the derivative of that

OpenStudy (amistre64):

i got no idea how to make the funny d so thisll be: \[grad(r^n) = grad(x^2+y^2+z^2)^{n/2}\]\[\hspace{5em}=\frac{n}{2}(x^2+y^2+z^2)^{(n-1)/2}*2x\frac{df}{dx},\frac{n}{2}(x^2+y^2+z^2)^{(n-1)/2}*2y\frac{df}{dy}\]\[\hspace{10em},\frac{n}{2}(x^2+y^2+z^2)^{(n-1)/2}*2z\frac{df}{dz}\]

OpenStudy (amistre64):

\[(x(x^2+y^2+z^2)^{(n-1)/2}\frac{df}{dx})\ i+...\]\[(y(x^2+y^2+z^2)^{(n-1)/2}\frac{df}{dy})\ j+...\]\[(z(x^2+y^2+z^2)^{(n-1)/2}\frac{df}{dz})\ k\] maybe

OpenStudy (amistre64):

forgot my ns :) xn yn and zn is what x y z are spose to be

OpenStudy (amistre64):

r^n R looks to be a scalar; then gradiented

OpenStudy (anonymous):

@.@

OpenStudy (anonymous):

oo

OpenStudy (amistre64):

i got no idea how to make this make sense tho lol

OpenStudy (amistre64):

the upside down triangle is notation for gradient; which is just a derivative notation as well

OpenStudy (amistre64):

and i prolly got some spurious looking dfs in there in hindsight

OpenStudy (anonymous):

\[r ^{n}=\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2}, R=<x,y,z>\] the scalar product between them is it \[<x\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2}, y\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2},z\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2}>\]? if yes, then the del of it is it\[<nx ^{2}\left( x ^{2}+y ^{2}+z ^{2} \right)^{(n/2)-1}, ny ^{2}\left( x ^{2}+y ^{2}+z ^{2} \right)^{(n/2)-1},nz ^{2}\left( x ^{2}+y ^{2}+z ^{2} \right)^{(n/2)-1}>\]?

OpenStudy (anonymous):

amistre, sry to bother you, can check whether my answer is correct, given if part a is correct?

OpenStudy (amistre64):

\[<x\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2}, y\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2},z\left( x ^{2}+y ^{2}+z ^{2} \right)^{n/2}>\] the scalar part looks good, assuming we are onthe right track. one issue is the you have\[\nabla . \vec {rR}\] which looks more like its asking for the dot product of the gradient and the rR .. unless thats a typo

OpenStudy (amistre64):

From: http://www.mathrec.org/vector.html The divergence operator is normally written as the dot product between a nabla and a vector function. When fonts do not allow, the divergence is usually written as div f. Verbally, the divergence of f is often refered to as "del dot f". The divergence is a scalar quantity. A serious discussion of the physical significance of the divergence is beyond the scope of this page, but if you think of the vector field as a fluid flow or a current, the divergence shows where fluid or charge is created or destroyed. The Cartesian form is given by: div f = (∂/∂x, ∂/∂y, ∂/∂z) · (fx, fy, fz) = ∂fx/∂x + ∂fy/∂y + ∂fz/∂z

OpenStudy (amistre64):

as long as we have defined r^n R correctly; then we simply add up the partial derivatives of it is what it looks like to me. \[\vec A = \vec {r^nR}\] \[\nabla.\vec A = A_x+A_y+A_z \] where those are the partial derivatives of A with respect to x y and z